# TokBench: Evaluating Your Visual Tokenizer before Visual Generation

### v1

@@ A.1 Tokenizer Selection
Resize 1.0 1.0 1.0 1.0 0.01 0.00 inf 1.00
Discrete VQGAN 0.10 0.19 0.47 0.25 4.27 0.13 23.98 0.72
Chameleon 0.19 0.30 0.58 0.36 3.63 0.11 24.54 0.73
LlamaGen 0.15 0.27 0.57 0.33 3.45 0.10 24.77 0.76
VAR 0.23 0.34 0.62 0.40 6.21 0.12 23.67 0.77
OmniTokenizer 0.40 0.60 0.82 0.61 4.63 0.15 16.00 0.71
TokenFlow 0.14 0.26 0.55 0.31 4.63 0.10 25.00 0.76
UniTok 0.33 0.50 0.76 0.53 3.78 0.07 26.54 0.84
O-MAGVIT2 0.20 0.34 0.66 0.40 3.62 0.09 26.01 0.79
O-MAGVIT2(pretrain) 0.21 0.34 0.65 0.40 3.48 0.09 26.12 0.79
LlamaGen(F8) 0.49 0.66 0.87 0.67 2.11 0.04 28.89 0.87
O-MAGVIT2(F8) 0.57 0.74 0.91 0.74 2.11 0.05 29.92 0.89
Continuous DC-AE 0.27 0.45 0.78 0.50 1.44 0.05 29.48 0.84
VA-VAE 0.49 0.68 0.89 0.69 2.38 0.04 30.69 0.88
SD-XL 0.50 0.69 0.91 0.70 1.25 0.03 31.39 0.89
SD-3.5 0.86 0.94 0.99 0.93 0.42 0.01 33.07 0.96
FLUX.1-dev 0.92 0.97 0.99 0.96 0.24 0.01 33.61 0.97

### v2

@@ A.1 Tokenizer Selection
Resize 1×1$\times$1 × 1.0 1.0 1.0 1.0 0.01 0.00 inf 1.00
Discrete VQGAN 16×16$\times$16 × 0.10 0.19 0.47 0.25 4.27 0.13 23.98 0.72
Chameleon 16×16$\times$16 × 0.19 0.30 0.58 0.36 3.63 0.11 24.54 0.73
LlamaGen 16×16$\times$16 × 0.15 0.27 0.57 0.33 3.45 0.10 24.77 0.76
VAR 16×16$\times$16 × 0.23 0.34 0.62 0.40 6.21 0.12 23.67 0.77
TokenFlow 16×16$\times$16 × 0.14 0.26 0.55 0.31 4.63 0.10 25.00 0.76
O-MAGVIT2 16×16$\times$16 × 0.20 0.34 0.66 0.40 3.62 0.09 26.01 0.79
O-MAGVIT2(pretrain) 16×16$\times$16 × 0.21 0.34 0.65 0.40 3.48 0.09 26.12 0.79
UniTok 16×16$\times$16 × 0.33 0.50 0.76 0.53 3.78 0.07 26.54 0.84
OmniTokenizer 8×8$\times$8 × 0.40 0.60 0.82 0.61 4.63 0.15 16.00 0.71
LlamaGen 8×8$\times$8 × 0.49 0.66 0.87 0.67 2.11 0.04 28.89 0.87
O-MAGVIT2 8×8$\times$8 × 0.57 0.74 0.91 0.74 2.11 0.05 29.92 0.89
Continuous DC-AE 32×32$\times$32 × 0.27 0.45 0.78 0.50 1.44 0.05 29.48 0.84
VA-VAE 16×16$\times$16 × 0.49 0.68 0.89 0.69 2.38 0.04 30.69 0.88
SD-XL 8×8$\times$8 × 0.50 0.69 0.91 0.70 1.25 0.03 31.39 0.89
SD-3.5 8×8$\times$8 × 0.86 0.94 0.99 0.93 0.42 0.01 33.07 0.96
FLUX.1-dev 8×8$\times$8 × 0.92 0.97 0.99 0.96 0.24 0.01 33.61 0.97